OpenStudy (anonymous):

EXAM TOMORROW already have work done, just need explanation. Solve (2x-1)/(x+3)

OpenStudy (anonymous):

then (2x-1)/(x+3)-1(x+3)/(x+3)</0

OpenStudy (anonymous):

I dont understand where the 1 went?

OpenStudy (anonymous):

and i dont get why you subtract x+3

OpenStudy (jdoe0001):

hmmm

OpenStudy (jdoe0001):

\(\large \bf \cfrac{2x-1}{x+3}\le 1?\)

OpenStudy (anonymous):

then he put (2x-1-x-3)/(x+3)</0

OpenStudy (anonymous):

yes

OpenStudy (jdoe0001):

well.... think about it this way \(\bf \cfrac{a}{b}\le 1\) if you subtract 1 from both sides.. you'd end up with \(\bf \cfrac{a}{b}-1\le \cancel{ 1-1 }\implies \cfrac{a}{b}-1\le 0\)

OpenStudy (jdoe0001):

you'd work with an inequality the same way you'd simplify a linear equation so... there's an assumption you do know how to simplify linear equations but the procedure is pretty much the same.... with one exception, but that doesn't apply here

OpenStudy (jdoe0001):

the easier way to go about it in this case, is just a cross-multiplication though

OpenStudy (anonymous):

why did he subtract x+3 though? and not on the other side?

OpenStudy (anonymous):

and if its subtracted on the bottom too, why does x+3 not change on the bottom?

OpenStudy (anonymous):

sorry I'm just really confused

OpenStudy (jdoe0001):

well.... that looks odd.... but I see what he did gimme one sec

OpenStudy (jdoe0001):

\(\bf \cfrac{2x-1}{x+3}\le 1\implies \cfrac{2x-1}{x+3}-1\le 0\qquad \begin{cases} \frac{\cancel{ same }}{\cancel{ same }}\to 1 \\ \quad \\ \frac{\cancel{ whatever }}{\cancel{ whatever }}\to 1\qquad thus \\ \quad \\ \frac{\cancel{ (x+3) }}{\cancel{ (x+3) }}\to 1\leftarrow \frac{(x+3) }{(x+3) } \end{cases} \\ \quad \\ \cfrac{2x-1}{x+3}-1\le 0\implies \cfrac{2x-1}{x+3}-\cfrac{(x+3) }{(x+3) }\le 0\implies \cfrac{(2x-1)\ -\ (x+3)}{x+3}\le 0 \\ \quad \\ \cfrac{2x-1-x-3}{x+3}\le 0\)

OpenStudy (jdoe0001):

\(\bf \cfrac{\cancel{ goose }}{\cancel{ goose }}=1\qquad \cfrac{\cancel{ fish }}{\cancel{ fish }}=1\qquad \cfrac{\cancel{ whatever }}{\cancel{ whatever }}=1\) thus

OpenStudy (jdoe0001):

thus \(\bf \cfrac{\cancel{ (x+3) }}{\cancel{ (x+3) }}\to 1\leftarrow \cfrac{(x+3) }{(x+3) }\)

OpenStudy (anonymous):

oh okay I see